# Offline Dedicated Quantum Attacks on Block Ciphers Based on Two Parallel Permutation-Based Pseudorandom Functions

### v2

@@ 2.5 Dedicated Quantum Attack on XOR-Type Function
The original space $\{0,1\}^{n}$ is partitioned into a $t$-dimensional linear subspace $\mathcal{L}=\{u:u=0^{n-t}\|*\}$ to construct a new function:
$F^{\mathcal{L}}(k,x^{l})=\bigoplus_{u\in\mathcal{L}}f(k,x^{l}\|0^{t}\oplus u).$
Using truncated technique, the parameters are decomposed as: $\alpha_{k}=\alpha_{k}^{l}\|\alpha_{k}^{r}$, $x=x^{l}\|x^{r}$, $s=s^{l}\|s^{r}$, $k=k^{l}\|k^{r}$, where $|\alpha_{k}^{l}|=|x^{l}|=|s^{l}|=|k^{l}|=n-t$ and $|\alpha_{k}^{r}|=|x^{r}|=|s^{r}|=|k^{r}|=t$. The resulting function $F^{\mathcal{L}}$ exhibits a truncated ntn-t-bit period $s^{l}$ when $\alpha_{k}^{l}=\alpha_{k_{0}}^{l}$, as stated in Property 1 [23].
Property 1 (Adapted from [23]).
Let $\mathcal{L}=\{u:u=0^{n-t}\|*\}$ and $f(k,x)$ be an XOR-type function such that $f(k,\cdot)$ has a non-trivial period $s$ when $k=k_{0}$. Then $F^{\mathcal{L}}(k,x^{l})=\bigoplus_{u\in\mathcal{L}}f(k,x^{l}\|0^{t}\oplus u)$ has an (nt)(n-t)-bit period $s^{l}$ for $\alpha^{l}_{k}=\alpha^{l}_{k_{0}}$.
In particular, if $f(k,x)=g_{1}(x)\oplus g_{2}(k\oplus x)$, then $F^{\mathcal{L}}(k^{l},x^{l})=\bigoplus_{u\in\mathcal{L}}f((k^{l}\oplus x^{l})\|0^{t}\oplus u)$ has an (nt)(n-t)-bit period $s^{l}$ for $k^{l}=k_{0}^{l}$.
The dedicated quantum attack on XOR-type function was proposed by Shi et al. [23], as presented in Theorem 3 and Algorithm 3.
Algorithm 3 Dedicated Attack on XOR-Type Function Using Truncated Technique [23]
Require: $\kappa,n,m,c^{\prime},t,r$, $f:\{0,1\}^{\kappa}\times\{0,1\}^{n}\to\{0,1\}^{m}$ of XOR-type, to which we have quantum oracle access;
@@
10:Compute the period $s^{l}$ by solving the equation system $s^{l}\cdot v_{i}=1$, $i=1,2,\ldots,c^{\prime}$;
Theorem 3.
(Adapted from [23]) Let $f:\{0,1\}^{\kappa}\times\{0,1\}^{n}\to\{0,1\}^{m}$ be a function of XOR-type, then there exists a quantum attack that uses $2^{(\kappa-t)/2}$ iterations with $c^{\prime}$ quantum queries per iteration. The probability of falling below a success rate of
@@
is lower than $2^{n+\kappa-2t-\frac{2^{n-t}}{4c^{\prime}}}$, where $c^{\prime}=n+\kappa-2t+\tau+1$.
Through the internal parallelization strategy, the attack needs $2^{(\kappa-t)/2}$ Grover iterations, each requiring $c^{\prime}=2n-2t+\tau+1$ Simon subroutine queries and $2^{t}$ quantum queries to $f$, thus this attack requires $2^{(\kappa+t)/2}$ quantum queries to the encryption oracle. The high (nt)(n-t)-bit values $\alpha_{k_{0}}^{l}$ (or $k_{0}^{l}$) and the truncated period $s^{l}$ are recovered first. The remaining $t$-bit $k^{r}$ is then recovered using Grover-meets-Simon algorithm applied to the function $f^{\prime}(k^{r},x)=f^{\prime}(k^{r},x\oplus s)$, where $f^{\prime}(k^{r},x)=f(k^{l}_{0}\|k^{r},x)$. Note that this part requires only $2^{t/2}$ iterations. This approach demonstrates significant improvements over general parallelization in terms of circuit depth, width, and gate count.
3 Offline Dedicated Quantum Attack on $p$-XOR-Type Function
In this section, we systematically explore offline dedicated quantum attack on block ciphers based on $p$-XOR-type function. Beginning with a new instance based on XOR-type function for online attack, we illustrate the obstacle faced when converting online attacks into offline ones. Then, we present other instantiations of TPP-PRFs and provide an offline attack for them, as these instances can be constructed as decoupled XOR-type function, i.e., $p$-XOR-type function.
3.1 A New Instance for Online Attack and Obstacle to Converting to Offline Attack
In this subsection, we find a BBB MAC, called PolyMAC, which can be attacked using quantum queries in Ref. [23]. Subsequently, we analyze the obstacle that prevents this attack to offline setting.
3.1.1 Quantum attack on PolyMAC for online query
PolyMAC scheme [15] is a Double-block Hash-then-Sum construction based on polynomial evaluation in 2020. It uses two hashing keys $k_{1},k_{3}\in\{0,1\}^{n}$ and two encryption keys $k_{2},k_{4}\in\{0,1\}^{m}$. We consider the case with two-block messages as follows:
@@ 3.1.1 Quantum attack on PolyMAC for online query
Next, we briefly describe the construction of an XOR-type function based on PolyMAC. Let $\beta_{0}$, $\beta_{1}$ be two fixed strings in $\{0,1\}^{n}$, $\beta_{0}\neq\beta_{1}$, and $\alpha_{k}\in\{0,1\}^{n}$, we define the function as
@@
$(\alpha_{i},x)$ $\mapsto\text{PolyMAC}(\beta_{0},x)\oplus\text{PolyMAC}\left(\beta_{1},x\oplus\alpha_{i}\right).$
It is easy to verify that the period $s$ is $(k_{1}\oplus k_{3})(\beta_{0}\oplus\beta_{1})$ when $\alpha_{i}=k_{1}(\beta_{0}\oplus\beta_{1})$ or $\alpha_{i}=k_{3}(\beta_{0}\oplus\beta_{1})$. From Eq. (4), we observe that the secret state $\alpha_{i}$ is XORed with $x$ in the second term $\text{PolyMAC}(\beta_{1},x\oplus\alpha_{i})$ of $f(\alpha_{i},x)$, which exhibits the XOR property.
Let $\mathcal{L}=\{u:u=0^{n-t}\|*\}$, $\alpha_{i}=\alpha_{i}^{l}\|\alpha_{i}^{r}$, $x=x^{l}\|x^{r}$, $|\alpha_{i}^{l}|=|x^{l}|=n-t$ and $|\alpha_{i}^{r}|=|x^{r}|=t$. Now, we propose a new function $F^{\mathcal{L}}$ with truncated input.
@@
According to Algorithm 3, we can recover the high (nt)(n-t) bits of the period $s^{l}=(k_{1}\oplus k_{3})^{l}(\beta_{0}\oplus\beta_{1})^{l}$ when $\alpha_{i}^{l}=k_{1}^{l}(\beta_{0}\oplus\beta_{1})^{l}$ or $k_{3}^{l}(\beta_{0}\oplus\beta_{1})^{l}$ with truncated technique. Consider $\alpha_{i}^{l}=k_{1}^{l}(\beta_{0}\oplus\beta_{1})^{l}$ as an example. After that, we can define $f^{\prime}(\alpha_{i}^{r},x)=f(k_{1}^{l}(\beta_{0}\oplus\beta_{1})^{l}\|\alpha_{i}^{r},x)$. Obviously, $f^{\prime}(\alpha_{i}^{r},x)=f^{\prime}(\alpha_{i}^{r},x\oplus s)$ when $\alpha_{i}^{r}=k_{1}^{r}(\beta_{0}\oplus\beta_{1})^{r}$, it requires $O(2^{t/2})$ quantum queries to $f$.
3.1.2 Obstacle to Converting to Offline Attack
@@
Although Section 3.1.1 identified a PolyMAC that can be attacked using dedicated attack proposed in Ref. [23], the difficulty of this attack lies in the requirement for quantum queries to $f$.
@@
The obstacle that prevents the conversion of a dedicated quantum attack from an online to an offline setting is the tightly coupled property of XOR-type function $f(i,x)=g_{1}(x)\oplus g_{2}(x\oplus\alpha_{i})$ based on block ciphers. The core of the issue lies in the functional coupling between the two components, $g_{1}$ and $g_{2}$, which requires queries to the encryption oracle. This construction of $f(i,x)$ binds the entire query process to interactive online queries, thereby completely precluding any possibility of offline computation.
@@
Specifically, consider the construction of XOR-type function based on PolyMAC given by Eq. (4), the first term $g_{1}(x)=\text{PolyMAC}(\beta_{0},x)$ and the second term $g_{2}(x\oplus\alpha_{i})=\text{PolyMAC}(\beta_{1},x\oplus\alpha_{i})$ of $f(\alpha_{i},x)$ both require quantum queries to the encryption oracle. We refer to the above structure as coupled XOR-type function. Similarly, XOR-type function constructed by several BBB MACs, which is proposed by Shi et al. [23], are all coupled XOR-type function.
@@
Therefore, we introduce the decoupled XOR-type function to provide the offline attack.
@@
decoupled XOR-type function. Let $f(i,x)$ be an XOR-type function in the following form $f:\{0,1\}^{\kappa}\times\{0,1\}^{n}\rightarrow\{0,1\}^{m}$, that is, $f(i,x)=g_{1}(x)\oplus g_{2}(x\oplus\alpha_{i})$, where $\alpha_{i}$ is the $n$-bit secret state determined by $i$. If the second term $g_{2}(x\oplus\alpha_{i})$ of $f(i,x)$ is a public function, which enables offline access, we call $f(i,x)$ is decoupled XOR-type function. To highlight its decoupled property, we denote it as XOR-type function with a public function ($p$-XOR-type function) and change its expression: $f(i,x)=g_{1}(x)\oplus p(x\oplus\alpha_{i})$, where $p(x\oplus\alpha_{i})$ represents a public function.
@@
To facilitate discussion, we consider the case where $i$ is XORed on $x$, we introduce the following notation. Let $\mathcal{L}=\{u:u=0^{n-t}||*\}$, the new periodic function $F^{\mathcal{L}}(i^{l},x^{l})=\bigoplus\limits_{u\in\mathcal{L}}f(i^{l},x^{l}||0^{t}\oplus u)=G_{1}^{\mathcal{L}}(x^{l})\oplus P^{\mathcal{L}}(i^{l},x^{l})$, where $G_{1}^{\mathcal{L}}(x^{l})=\bigoplus\limits_{u\in\mathcal{L}}g_{1}(x^{l}\|0^{t}\oplus u)$, $P^{\mathcal{L}}$ consists of the public function $p(i,x)$, i.e., $P^{\mathcal{L}}(i^{l},x^{l})=\bigoplus\limits_{u\in\mathcal{L}}p((x^{l}\oplus i^{l})\|0^{t}\oplus u)$. It is obvious that $p$-XOR-type function is a special class of XOR-type function. Thus, Property 1 and Algorithm 3 proposed by Shi et al. [23] holds for $p$-XOR-type function.
3.2 Several Other Instances for Offline Attack
@@
Given that the decoupling property of XOR-type function is crucial for enabling offline attack, we next explore a class of structure, TPP-PRFs, that inherently possess this characteristic, and present $p$-XOR-type function constructed from them.
3.2.1 Instantiation of TPP-PRFs
Let us show that $p$-XOR-type function constructed by instantiations of TPP-PRFs, and construct new periodic functions.
@@ 3.2.1 Instantiation of TPP-PRFs
TPP-PRFs [7] TPP-PRFs with $n$-bit input $x$, which are constructed with two parallel permutations $P_{1}$ and $P_{2}$, present $2n/3$ bits of security in the classical setting, defined as follows.
@@
$x$ $\mapsto l_{33}P_{1}(l_{13}(x)\oplus l_{14}(k_{1}))\oplus l_{34}P_{2}(l_{23}(x)\oplus l_{24}(k_{2}))\oplus e(x)\oplus C$
where $e(x)=l_{31}l_{11}(x)\oplus l_{32}l_{21}(x)$, the constant term $C=l_{31}l_{12}(k_{1})\oplus l_{32}l_{22}(k_{2})\oplus l_{35}(k_{3})\oplus l_{36}(k_{4})$, and $l_{ij}\neq 0$ for $i=1,2,3$, $j=3,4$.
Based on TPP-PRFs, we can construct a hidden periodic function that is a $p$-XOR-type function.
@@
$=l_{33}P_{1}(l_{13}(x\oplus l_{13}^{-1}l_{14}(k_{1}))\oplus l_{34}P_{2}(l_{23}(x\oplus l_{23}^{-1}l_{24}(k_{2}))$
$\quad\oplus l_{33}P_{1}\left(l_{13}(x)\right)\oplus l_{34}P_{2}\left(l_{23}(x\oplus i)\right)\oplus C,$
where $g_{1}(x)=g(x)\oplus l_{33}P_{1}\left(l_{13}(x)\right)\oplus e(x)$, and the public function $p:\{0,1\}^{n}\times\{0,1\}^{n}\rightarrow\{0,1\}^{n}$, that is, $p(i,x)=l_{34}P_{2}\left(l_{23}(x\oplus i)\right)$. It can be verified that $f(i,x)=f(i,x\oplus l_{13}^{-1}l_{14}(k_{1}))$ when $i=l_{23}^{-1}l_{24}(k_{2})$ or $i=l_{23}^{-1}l_{24}(k_{2})\oplus l_{23}l_{13}^{-1}l_{14}(k_{1})$ in Ref. [31]. Since $i$ is XORed with $x$ in the public function $p(i,x)=l_{34}P_{2}(l_{23}(x\oplus i)$ of $f(i,x)$, the hidden periodic function $f(i,x)$ is a $p$-XOR-type function.
For $p$-XOR-type function constructed by TPP-PRFs described above, we can apply truncation techniques to obtain a new periodic function denoted as $F^{\mathcal{L}}$.
Let $\mathcal{L}=\{u:u=0^{n-t}\|*\}$, the (nt)(n-t)-bit key $k_{1}^{l}$ is decomposed into two parts: $k_{1}^{l_{1}}$ with $p$ bits and $k_{1}^{l_{2}}$ with (ntp)(n-t-p)-bits for $0\leq p\leq n-t$, where $t$ is a tunable positive integer truncation parameter. We define $G_{1}^{\mathcal{L}}:\{0,1\}^{p}\rightarrow\{0,1\}^{n}$ and the public function $P^{\mathcal{L}}:\{0,1\}^{2n-2t-p}\times\{0,1\}^{p}\rightarrow\{0,1\}^{n}$ as
$G_{1}^{\mathcal{L}}(x^{l_{1}})$ $=\bigoplus\limits_{u\in\mathcal{L}}g(x^{l_{1}}\|0^{n-t-p}\|0^{t}\oplus u)\oplus e(x^{l_{1}}\|0^{n-t-p}),$
$P^{\mathcal{L}}(i^{l}\|j^{l_{2}},x^{l_{1}})$ $=\bigoplus\limits_{u\in\mathcal{L}}l_{33}P_{1}\left(l_{13}(x^{l_{1}}\|j^{l_{2}})\right)\oplus l_{34}P_{2}\left(l_{23}\left((x^{l_{1}}\|0^{n-t-p})\oplus i\right)\right).$
Now, we construct the new periodic function $F^{\mathcal{L}}:\{0,1\}^{2n-2t-p}\times\{0,1\}^{p}\rightarrow\{0,1\}^{n}$ as follows,
$F^{\mathcal{L}}(i^{l}\|j^{l_{2}},x^{l_{1}})=$ $G_{1}^{\mathcal{L}}(x^{l_{1}})\oplus P^{\mathcal{L}}(i^{l}\|j^{l_{2}},x^{l_{1}})$ (8)
$=$ $\bigoplus\limits_{u\in\mathcal{L}}g(x^{l_{1}}\|0^{n-t-p}\|0^{t}\oplus u)\oplus l_{33}P_{1}\left(l_{13}(x^{l_{1}}\|j^{l_{2}})\right)$
$\oplus l_{34}P_{2}\left(l_{23}\left((x^{l_{1}}\|0^{n-t-p})\oplus i\right)\right)\oplus e(x^{l_{1}}\|0^{n-t-p}),$
@@
In particular, this function has the period $s^{l_{1}}=l_{13}^{-1}l_{14}k_{1}^{l_{1}}$ if and only if $i^{l}\|j^{l_{2}}=l_{23}^{-1}l_{24}k_{2}^{l}\|l_{13}^{-1}l_{14}k_{1}^{l_{2}}$ (or $i^{l}\|j^{l_{2}}=l_{23}^{-1}l_{24}k_{2}^{l}\oplus l_{23}l_{13}^{-1}l_{14}k_{1}^{l}\|l_{13}^{-1}l_{14}k_{1}^{l_{2}}$). As an example, we take $i^{l}\|j^{l_{2}}=l_{23}^{-1}l_{24}k_{2}^{l}\|l_{13}^{-1}l_{14}k_{1}^{l_{2}}$ to show that

### v3

@@ II-D Dedicated Quantum Attack on Constructions Based on XOR-Type Function
A new function $F^{\mathcal{L}}$ is constructed by summing the value of $f$ over the $t$-dimensional linear subspace $\mathcal{L}=\{u:u=0^{n-t}\|*\}$. The function $F^{\mathcal{L}}$ is written as
$F^{\mathcal{L}}(k,x^{l})=\bigoplus_{u\in\mathcal{L}}f(k,(x^{l}\|0^{t})\oplus u).$
Using the truncated technique, the $n$-bit parameters are decomposed into (nt)(n-t)-bit parts and $t$-bit parts: $X=X^{l}\|X^{r}$, where $X\in\{\alpha_{k},x,s\}$, $|X^{l}|=n-t$, and $|X^{r}|=t$. The function $F^{\mathcal{L}}$ exhibits an (nt)(n-t)-bit truncated period $s^{l}$ when $\alpha_{k}^{l}=\alpha_{k_{0}}^{l}$, as detailed in Property 1 [shi2024dedicated].
Property 1
(Adapted from [shi2024dedicated]) Let $\mathcal{L}=\{u:u=0^{n-t}\|*\}$ and $f(k,x)$ be an XOR-type function such that $f(k,\cdot)$ has a non-trivial period $s$ when $k=k_{0}$. Then $F^{\mathcal{L}}(k,x^{l})=\bigoplus_{u\in\mathcal{L}}f(k,x^{l}\|0^{t}\oplus u)$ has an (nt)(n-t)-bit period $s^{l}$ for $\alpha^{l}_{k}=\alpha^{l}_{k_{0}}$.
In particular, if $f(k,x)=g_{1}(x)\oplus g_{2}(k\oplus x)$, then $F^{\mathcal{L}}(k^{l},x^{l})=\bigoplus_{u\in\mathcal{L}}f(k^{l},(x^{l}\|0^{t})\oplus u)$ has an (nt)(n-t)-bit period $s^{l}$ for $k^{l}=k_{0}^{l}$.
The dedicated quantum attack on constructions based on XOR-type functions using truncated technique was proposed by Shi et al. [shi2024dedicated], and is summarized in Algorithm 3 and Theorem 3 based on the original description.
Algorithm 3 Dedicated Attack on Constructions Based on XOR-Type Function Using Truncated Technique [shi2024dedicated]
Require: $\kappa,n,m,c^{\prime},t,r$. Let $f:\{0,1\}^{\kappa}\times\{0,1\}^{n}\to\{0,1\}^{m}$ be an XOR-type function to which we have quantum oracle access;
@@
10:Compute the period $s^{l}$ by solving the equation system $s^{l}\cdot v_{i}=0$, $i=1,2,\ldots,c^{\prime}$;
Theorem 3
(Adapted from [shi2024dedicated]) Let $f:\{0,1\}^{\kappa}\times\{0,1\}^{n}\to\{0,1\}^{m}$ be an XOR-type function, there exists a quantum attack that uses $2^{(\kappa-t)/2}$ iterations with $c^{\prime}$ quantum queries per iteration. The probability of falling below a success rate of
@@
is lower than $2^{n+\kappa-2t-\frac{2^{n-t}}{4c^{\prime}}}$, where $0<t<n$, $c^{\prime}=n+\kappa-2t+\tau+1$.
The quantum complexity of this attack is analyzed as follows. A larger truncation parameter $t$ results in fewer iterations of $2^{(\kappa-t)/2}$ but yields a shorter recovered key of (nt)(n-t)-bit, which consists of $\alpha_{k_{0}}^{l}$ and the period $s^{l}$. Each iteration requires $c^{\prime}$ queries to the function $F^{\mathcal{L}}$. Since $F^{\mathcal{L}}$ is a sum over $2^{t}$ parallel evaluations of $f$, one query to $F^{\mathcal{L}}$ is counted as $2^{t}$ quantum queries to the underlying function $f$. Consequently, the total quantum query complexity to the encryption oracle is $O(2^{(\kappa-t)/2}\cdot 2^{t}\cdot(n-t))=O(2^{(\kappa+t)/2}\cdot(n-t))$.
The remaining $t$-bit $k^{r}$ is then recovered using the Grover-meets-Simon algorithm on the function $f^{\prime}(k^{r},x)$, defined as $f^{\prime}(k^{r},x)=f(k^{l}_{0}\|k^{r},x)$. This function exhibits a period $s$ when the key $k^{l}_{0}\|k^{r}$ is correct. Crucially, this part requires only $O(2^{t/2})$ iterations.
III Offline Dedicated Quantum Attack on Block Ciphers by Constructing Decoupled XOR-Type Function
In this section, we propose an offline dedicated quantum attack on block ciphers by constructing decoupled XOR-type functions. We begin by presenting a new instance based on an XOR-type function for an online attack, and then we illustrate the obstacle encountered when converting it into an offline attack. Subsequently, we propose an offline attack on other instantiations of TPP-PRFs, which can be constructed as decoupled XOR-type functions.
III-A A New Instance for Online Attack and Obstacle to Offline Conversion
In this subsection, we identify PolyMAC, a BBB MAC that is subject to the dedicated quantum attack in Ref. [shi2024dedicated]. Subsequently, we analyze the obstacle that prevents converting this attack to an offline setting.
III-A1 Quantum Attack on PolyMAC with Online Queries
PolyMAC [kim2020tight]. PolyMAC scheme is a Double-block Hash-then-Sum construction based on polynomial evaluation, proposed in 2020. It uses two hashing keys $k_{1},k_{3}\in\{0,1\}^{n}$ and two encryption keys $k_{2},k_{4}\in\{0,1\}^{m}$. We consider the case with the two-block message as follows:
@@ III-A1 Quantum Attack on PolyMAC with Online Queries
Next, we describe the construction of an XOR-type function based on PolyMAC. Let $\beta_{0}$, $\beta_{1}$ be two fixed strings in $\{0,1\}^{n}$, $\beta_{0}\neq\beta_{1}$, and $\alpha_{k}\in\{0,1\}^{n}$, we define the function as
@@
$(\alpha_{i},x)$ $\mapsto\text{PolyMAC}(\beta_{0},x)$
$\oplus\text{PolyMAC}\left(\beta_{1},x\oplus\alpha_{i}\right).$
The period $s$ is $(k_{1}\oplus k_{3})(\beta_{0}\oplus\beta_{1})$ when $\alpha_{i}=k_{1}(\beta_{0}\oplus\beta_{1})$ or $\alpha_{i}=k_{3}(\beta_{0}\oplus\beta_{1})$. From Eq. (4), the secret state $\alpha_{i}$ is XORed with $x$ in the second term $\text{PolyMAC}(\beta_{1},x\oplus\alpha_{i})$ of $f(\alpha_{i},x)$.
Now, we propose a new function $F^{\mathcal{L}}$ using truncated input. Let $\mathcal{L}=\{u:u=0^{n-t}\|*\}$, $\alpha_{i}=\alpha_{i}^{l}\|\alpha_{i}^{r}$, $x=x^{l}\|x^{r}$, $|\alpha_{i}^{l}|=|x^{l}|=n-t$, and $|\alpha_{i}^{r}|=|x^{r}|=t$.
@@
The function $F^{\mathcal{L}}$ has a period $s^{l}=(k_{1}\oplus k_{3})^{l}(\beta_{0}\oplus\beta_{1})^{l}$ when $\alpha_{i}^{l}=k_{1}^{l}(\beta_{0}\oplus\beta_{1})^{l}$ or $k_{3}^{l}(\beta_{0}\oplus\beta_{1})^{l}$. By applying Algorithm 3, we can recover the high (nt)(n-t)-bits $s^{l}$ and $\alpha_{i}^{l}$.
@@
Taking $\alpha_{i}^{l}=k_{1}^{l}(\beta_{0}\oplus\beta_{1})^{l}$ as an example, we define $f^{\prime}(\alpha_{i}^{r},x)=f(k_{1}^{l}(\beta_{0}\oplus\beta_{1})^{l}\|\alpha_{i}^{r},x)$. When $\alpha_{i}^{r}=k_{1}^{r}(\beta_{0}\oplus\beta_{1})^{r}$, $f^{\prime}(\alpha_{i}^{r},x)=f^{\prime}(\alpha_{i}^{r},x\oplus s)$, which requires $O(2^{t/2})$ quantum queries to $f$.
III-A2 Obstacle to Offline Conversion
@@
Although Section III-A1 showed that PolyMAC is vulnerable to the attack proposed in Ref. [shi2024dedicated], this attack is limited by its requirement for quantum queries to the encryption oracle.
@@
The obstacle to converting this dedicated quantum attack from an online to an offline setting is the tightly coupled property of an XOR-type function $f(i,x)=g_{1}(x)\oplus g_{2}(x\oplus\alpha_{i})$. The issue is that both $g_{1}$ and $g_{2}$ require queries to the encryption oracle. This construction of $f(i,x)$ binds the entire query process to interactive online queries, thus precluding the possibility of offline computation.
@@
Specifically, consider the construction of an XOR-type function based on PolyMAC given by Eq. (4), both the first term $g_{1}(x)=\text{PolyMAC}(\beta_{0},x)$ and the second term $g_{2}(x\oplus\alpha_{i})=\text{PolyMAC}(\beta_{1},x\oplus\alpha_{i})$ of $f(\alpha_{i},x)$ require quantum queries to the encryption oracle. We refer to this as a coupled XOR-type function. Similarly, all XOR-type functions constructed by several BBB MACs, which are proposed by Shi et al. [shi2024dedicated], are coupled XOR-type functions.
@@
Therefore, we introduce a decoupled XOR-type function to enable an offline attack.
@@
Decoupled XOR-type function. Let $f:\{0,1\}^{\kappa}\times\{0,1\}^{n}\rightarrow\{0,1\}^{m}$ be an XOR-type function of the form $f(i,x)=g_{1}(x)\oplus g_{2}(x\oplus\alpha_{i})$, where $\alpha_{i}$ is the $n$-bit secret state determined by $i$. If the second term $g_{2}(x\oplus\alpha_{i})$ of $f(i,x)$ is a public function that enables offline computation, we call $f(i,x)$ a decoupled XOR-type function. To highlight its decoupled property, we denote it as an XOR-type function with a public function ($p$-XOR-type function) and express it as $f(i,x)=g_{1}(x)\oplus p(x\oplus\alpha_{i})$, where $p(x\oplus\alpha_{i})$ is a public function.
@@
To facilitate discussion, we introduce the following notation for the case where $i$ is XORed on $x$. Let $\mathcal{L}=\{u:u=0^{n-t}||*\}$. We define a new periodic function $F^{\mathcal{L}}(i^{l},x^{l})$ as follows: $F^{\mathcal{L}}(i^{l},x^{l})=\bigoplus\limits_{u\in\mathcal{L}}f(i^{l},x^{l}||0^{t}\oplus u)=G_{1}^{\mathcal{L}}(x^{l})\oplus P^{\mathcal{L}}(i^{l},x^{l})$, where $G_{1}^{\mathcal{L}}(x^{l})=\bigoplus\limits_{u\in\mathcal{L}}g_{1}(x^{l}\|0^{t}\oplus u)$. Here, $P^{\mathcal{L}}$ is defined based on the public function $p(i,x)$ as $P^{\mathcal{L}}(i^{l},x^{l})=\bigoplus\limits_{u\in\mathcal{L}}p((x^{l}\oplus i^{l})\|0^{t}\oplus u)$. The $p$-XOR-type function is a special class of the XOR-type function. Consequently, Property 1 and Algorithm 3 proposed by Shi et al. [shi2024dedicated] are also applicable to the $p$-XOR-type function.
III-B Other Instances for Offline Attack
@@
Given that the decoupled property of XOR-type functions is crucial for implementing an offline attack, we next explore TPP-PRFs, whose periodic functions naturally possess this property.
III-B1 Instantiation of TPP-PRFs
We show how to construct $p$-XOR-type functions from instantiations of TPP-PRFs.
@@ III-B1 Instantiation of TPP-PRFs
TPP-PRFs [chen2019build] TPP-PRFs, which use $n$-bit input $x$ and are constructed with two parallel permutations $P_{1}$ and $P_{2}$, achieve $2n/3$ bits of security in the classical setting, defined as follows.
@@
$x$ $\mapsto l_{33}P_{1}(l_{13}(x)\oplus l_{14}(k_{1}))$
$~~~~~\oplus l_{34}P_{2}(l_{23}(x)\oplus l_{24}(k_{2}))\oplus e(x)\oplus C$
where $e(x)=l_{31}l_{11}(x)\oplus l_{32}l_{21}(x)$, the constant term is $C=l_{31}l_{12}(k_{1})\oplus l_{32}l_{22}(k_{2})\oplus l_{35}(k_{3})\oplus l_{36}(k_{4})$, and $l_{ij}\neq 0$ for $i=1,2,3$, $j=3,4$.
Based on TPP-PRFs, we can construct the following function that is a $p$-XOR-type function.
@@
where $g_{1}(x)=g(x)\oplus l_{33}P_{1}\left(l_{13}(x)\right)\oplus e(x)$, and the public function $p:\{0,1\}^{n}\times\{0,1\}^{n}\rightarrow\{0,1\}^{n}$, that is, $p(i,x)=l_{34}P_{2}\left(l_{23}(x\oplus i)\right)$. It can be verified that $f(i,x)=f(i,x\oplus l_{13}^{-1}l_{14}(k_{1}))$ when $i=l_{23}^{-1}l_{24}(k_{2})$ or $i=l_{23}^{-1}l_{24}(k_{2})\oplus l_{23}l_{13}^{-1}l_{14}(k_{1})$ in Ref. [sun2025quantum]. Since $i$ is XORed with $x$ in the public function $p(i,x)=l_{34}P_{2}(l_{23}(x\oplus i)$ of $f(i,x)$, the function $f(i,x)$ is a $p$-XOR-type function.
For $p$-XOR-type functions constructed by TPP-PRFs, we can apply truncation techniques to obtain a new periodic function $F^{\mathcal{L}}$.
Let $\mathcal{L}=\{u:u=0^{n-t}\|*\}$. We decompose the (nt)(n-t)-bit key $k_{1}^{l}$ into two parts: $k_{1}^{l_{1}}$ with $p$ bits and $k_{1}^{l_{2}}$ with (ntp)(n-t-p)-bits, where $0\leq p\leq n-t$, and $t$ is a tunable positive integer truncation parameter. We define $G_{1}^{\mathcal{L}}:\{0,1\}^{p}\rightarrow\{0,1\}^{n}$ and the public function $P^{\mathcal{L}}:\{0,1\}^{2n-2t-p}\times\{0,1\}^{p}\rightarrow\{0,1\}^{n}$ as
$G_{1}^{\mathcal{L}}(x^{l_{1}})$ $=\bigoplus\limits_{u\in\mathcal{L}}g(x^{l_{1}}\|0^{n-t-p}\|0^{t}\oplus u)$
$~~~~\oplus e(x^{l_{1}}\|0^{n-t-p}\|0^{t}\oplus u),$
$P^{\mathcal{L}}(i^{l}\|j^{l_{2}},x^{l_{1}})$ $=\bigoplus\limits_{u\in\mathcal{L}}l_{33}P_{1}\left(l_{13}(x^{l_{1}}\|j^{l_{2}})\|0^{t}\oplus u\right)$
$~~~~\oplus l_{34}P_{2}\left(l_{23}\left((x^{l_{1}}\|0^{n-t-p})\oplus i^{l}\right)\|0^{t}\oplus u\right).$
We construct the new periodic function $F^{\mathcal{L}}:\{0,1\}^{n-t}\times\{0,1\}^{n-t-p}\times\{0,1\}^{p}\rightarrow\{0,1\}^{n}$ as follows:
$F^{\mathcal{L}}(i^{l}\|j^{l_{2}},x^{l_{1}})=$ $G_{1}^{\mathcal{L}}(x^{l_{1}})\oplus P^{\mathcal{L}}(i^{l}\|j^{l_{2}},x^{l_{1}}),$ (8)
@@
This function has the period $s^{l_{1}}=l_{13}^{-1}l_{14}k_{1}^{l_{1}}$ when $i^{l}\|j^{l_{2}}=l_{23}^{-1}l_{24}k_{2}^{l}\|l_{13}^{-1}l_{14}k_{1}^{l_{2}}$ (or $i^{l}\|j^{l_{2}}=l_{23}^{-1}l_{24}k_{2}^{l}\oplus l_{23}l_{13}^{-1}l_{14}k_{1}^{l}\|l_{13}^{-1}l_{14}k_{1}^{l_{2}}$). Take $i^{l}\|j^{l_{2}}=l_{23}^{-1}l_{24}k_{2}^{l}\|l_{13}^{-1}l_{14}k_{1}^{l_{2}}$ as an example